OpenStudy (cookiemonster18):

For the equation -4y=8x what is the constant or variations? A.-4 B.-2 C.1 D.2

OpenStudy (cookiemonster18):

I think it's A am I correct?

OpenStudy (yeval76):

Yea its A

OpenStudy (anonymous):

yea

OpenStudy (cookiemonster18):

Can you help with 3 more?

Nnesha (nnesha):

solve for y

Nnesha (nnesha):

equation for direct variation is `y=kx` equation for inverse variation is `y=k/x` where k is constant of variation

OpenStudy (cookiemonster18):

suppose y varies directly with x, and y=24 when x=8 what is the value of y when x=10? A.10 B.30 C.80 D.240 (I think C am I correct?) @Nnesha @ransondyara471 @yeval76

Nnesha (nnesha):

what did you get for first one ?

Nnesha (nnesha):

nd no direct variation y =kx plugin x =8 y=24 and then solve for k(constant of variation )

OpenStudy (cookiemonster18):

@Nnesha can you teach/show me how to get the answer ik you already did but I still don't understand how to solve to get the answer can you do it in steps please

Nnesha (nnesha):

for first one can you solve for y ?

Nnesha (nnesha):

how would you cancel out -4 from left side \[\large\rm -4y=8x\]?

Nnesha (nnesha):

no we don't need to add

Nnesha (nnesha):

it's -4 times y so to cancel out -4 do opposite of multiplication

Nnesha (nnesha):

opposite of multiplication is ??

OpenStudy (cookiemonster18):

division

Nnesha (nnesha):

right divide both sides by -4

OpenStudy (cookiemonster18):

-4/-4=1 8/-4=-2

Nnesha (nnesha):

correct!

OpenStudy (cookiemonster18):

so for the first one the correct answer is -2? and not -4?....

OpenStudy (cookiemonster18):

@Nnesha

Nnesha (nnesha):

yes

OpenStudy (cookiemonster18):

can you help me with the second one @Nnesha

Nnesha (nnesha):

what's the equation for `direct variation` ?

OpenStudy (cookiemonster18):

y=kx??

Nnesha (nnesha):

correct replace x and y for their values then solve for k

Nnesha (nnesha):

`y=24 when x=8 ` y=kx

OpenStudy (cookiemonster18):

so i'd divide 24 by??? k, but what is k. how do I find it? @Nnesha

Nnesha (nnesha):

first replace y with 24 and x with 8 to solve for k would you move 8 to the left side or 24 to the right side ?

OpenStudy (cookiemonster18):

8 to the right? 24/8=3 24/24=1

Nnesha (nnesha):

wayes 3 is the constant of variation

Nnesha (nnesha):

what do you mean by 24/24

Nnesha (nnesha):

\[\rm 24=8k\] divide both side by 8 you get k=3 now substitute k for 3 and x for 10 solve for y

OpenStudy (cookiemonster18):

I'm confuse. So itll be B. 30 for the answer

Nnesha (nnesha):

yes correct \[\rm y=kx\] k=3 so when x=10 y=? we should replace x with 10 and k with 3 then multiply

OpenStudy (cookiemonster18):

thank you. I had no idea how to solve problems like these I somewhat just learned how to do it a little. I'm not a pro but I some what know a little bit of what to do

Nnesha (nnesha):

yeah practice will help you to learn it:=))